Find the value of n
49/5 * n = 1

I'm not sure how to do this... i forgot. Please help in any way you can.

Found 2 solutions by edjones, Alan3354:
Answer by edjones(8007)

49/5 * n = 1
n=5/49

Answer by Alan3354(69443)

Find the value of n
49/5 * n = 1
--------
Assuming you mean (49/5)*n = 1
49/5 = 1/n
Invert
n = 5/49
-----------
Or,
(49/5)*n = 1
Multiply by 5
49n = 5
Divide by 49
n = 5/49